OpenStudy (lgg23):

Rewrite the expression as an allgebraic expresion in x: cos(arc cos x + arc sin x) = ?

OpenStudy (asnaseer):

 In the diagram above \(\cos(\alpha)=x\) and \(\sin(\beta)=x\). therefore:\[\cos(\arccos(x)+\arcsin(x))=\cos(\alpha+\beta)=\cos(90)=0\]